OpenStudy (anonymous):

5/9z^3 + 1/15z^2 5/9z^3 + 1/15z^2 @Mathematics

OpenStudy (anonymous):

\[\frac{1}{45} z^2 (25 z+3) \]

OpenStudy (anonymous):

where did you get the 45?

OpenStudy (anonymous):

\[\frac{5 z^3}{9}+\frac{z^2}{15}=\frac{15*5 z^3+9*z^2}{9*15}=\frac{75 z^3+9 z^2}{135}=\frac{3z^2 (25 z+3)}{3*45}= \frac{z^2 (3+25 z)}{45} \]